OpenStudy (anonymous):

Antonia works 40 hours per week at a eletronic warehouse. he earns 6.50 per hour plus a 3% commission on the total dollar value of the service contracts he sells. If Antonia's hourly rate was increased by 0.15 and his commission were raised to 5%. how much would he earn if he sold 4,000 dollars worth of service contracts for the week a)126.50 b)206.65 c)466.00 d)580.00

OpenStudy (anonymous):

6.50+0.15=6.65 -6.65 per hour 6.65(40) +0.05(4000) 266+200=?